OpenStudy (brsphysics):

a) A 35.5-m length of copper wire at 20.0°C has a radius of 0.27 mm. If a potential difference of 7.0 V is applied across the length of the wire, determine the current in the wire. (Use the values in this table.) b) If the wire is heated to 40.0°C while the 7.0-V potential difference is maintained, what is the resulting current in the wire?

OpenStudy (shamim):

do u know ohm's law

OpenStudy (shamim):

ohm's law is \[V=IR\]

OpenStudy (shamim):

u hv to use \[R=\frac{ \rho L }{ A }\]

OpenStudy (shamim):

u hv to use \[R _{\theta} =R _{0}(1+\alpha \theta)\]

OpenStudy (shamim):

anyway welcome to openstudy

OpenStudy (shamim):

response to get better help!!!